# In-2-4D: Inbetweening from Two Single-View Images to 4D Generation

### v2

@@ 3.2 Temporal Fragment Hierarchy
[…]
$\mathcal{K}={K_{(s)(1)},K_{(1)(2)},K_{(2)(3)},...,K_{(c)(e)}}$ (1)
@@
$\mathbf{V}=\{\mathcal{V}_{s(1f)},\mathcal{V}_{(1f)(2f)},...,\mathcal{V}_{((c-1)f)e}\},$ (2)
@@
3.3 Modelling Intra-Fragment Geometry
We lift individual video fragments to 4D by generating multi-view videos of the object. Existing video-to-4D methods [30, 41] use multi-view diffusion models [24] to synthesize multi-view videos by independently processing each frame. However, this approach ensures cross-view consistency but leads to temporally inconsistent geometry Fig Fig. 7. Moreover, due to sole reliance on multi-view video supervision, Gaussian splatting often produces flickering and texture variations [25] due to its high degrees of freedom per point and lack of motion constraints. We address these issues by generating temporally consistent multi-view videos and regularizing motion with rigid constraints within each fragment.
Learning canonical 3D. Similar to prior works, we first estimate a canonical Gaussian representation and then add motion to it from the multi-view videos. For each temporal fragment $\mathcal{V}{i}$, we designate the keyframe $K{(i)(j)}$ as the canonical reference and reconstruct its 3D structure via multi-view synthesis. Specifically, we employ the multiview diffusion model Era3D [18] to generate multi-view images from $K{(i)(j)}$, followed by 3DGS [13] for coarse geometry reconstruction. As each fragment is processed independently, parallel execution reduces computation time. The resulting coarse geometry provides an effective initialization for learning texture and geometry across the remaining temporal frames.
Dynamic 3D fragment generation. After learning 3D static Gaussians, we leverage motion priors from the video fragment to transform them into dynamic Gaussians. Since single-view videos cannot provide diverse observations of the scene from different viewpoints, we use multi-view videos. To promote temporal consistency, rather than generating multiple-view of the frames independently at each timestep, we propagate the self-attention features of the multi-view diffusion model [18] from the canonical frame across the entire frames of the fragment as follows:
@@ 3.3 Modelling Intra-Fragment Geometry
where ${z}_{c}$ is the multi-view latent of the canonical frame, $t$ is timestep of the video fragment, $\gamma$ is the blending weight and $d_{k}$ is the key dimension. With quasi-static motions in each video fragment, the generated multi-view videos have minimal variation in viewpoints, making it easier for the model to capture accurate and consistent geometry (Fig. 7). With the synthesized multi-view videos of the dynamic object, we optimize a 3D deformation field (denoted by $\Delta_{\boldsymbol{\Phi}_{i}}$) to enable free-viewpoint rendering. We chose Hexplanes [4] as our deformation field due to modeling efficiency. The deformation field predicts each Gaussian’s geometric offsets at a given timestamp relative to the mean canonical state (keyframe). For each timestamp $\tau$ of video and 3D Gaussian $p$, Hexplanes predict displacement, rotation, and scaling for the 3D gaussian points.
Figure 5: Effect of inbetweening on geometry. When the input states are significantly different, the 3D deformation module undergoes large movements (fast motion) leading to artifacts in novel views, whereas generating intermediate frames between the states (slow motion) enhances the geometry using smaller deformations.
@@
$\mathcal{L}_{\textrm{Ref}}=\frac{1}{\mathcal{T}}\sum_{\tau=1}^{\mathcal{T}}||f(\phi(S,\tau),o_{\textrm{Ref}})-I^{\tau}_{\textrm{Ref}}||^{2}_{2},$ (3)
where $I^{\tau}_{\textrm{Ref}}$ is the $\tau$-th frame, $o_{\textrm{Ref}}$ is the reference viewpoint, and $f$ is the rendering function. Dynamic Gaussians tend to move freely across regions of similar color [25] without constraints, causing flickering and floating artifacts that degrade 4D motion realism. As motion within each fragment is minimal, we enforce rigid assumptions on point movements relative to the canonical state by regularization:
$\mathcal{L}_{rigid}=||d(\mu^{c}_{i},\mu^{c}_{j})-d(\mu^{\tau}_{i},\mu^{\tau}_{j})||_{1}$ (4)
where $d(x,y)=||x-y||_{2}$ is the distance function, and $\mu$ denotes the Gaussian center of neighboring clusters $\mathcal{N}$. $\mu^{c}$ and $\mu^{\tau}$ represent Gaussian centers in the canonical and arbitrary timestep frames, respectively. This regularization permits non-rigid deformations (like bending) while minimizing local rigid distortions. In addition to this, we also use a random view at each timestep and apply foreground mask loss $\mathcal{L}_{mask}$, resulting in a total training objective:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{Ref}+\lambda_{2}\mathcal{L}_{mask}+\lambda_{3}\mathcal{L}_{rigid},$ (5)
where $\lambda_{1}$ and $\lambda_{2}$ are weights. In training, for each fragment $\mathcal{V}_{i}$, we first use $\mathcal{L}$ to supervise the static 3D Gaussian, then train the dynamic 4D Gaussian with all reference frames.
3.4 Cascaded 3D Motion Aggregation
@@
$\Delta=[\boldsymbol{\Delta_{\boldsymbol{\Phi}_{1}}},\boldsymbol{\Delta_{\boldsymbol{\Phi}_{2}}},...,\boldsymbol{\Delta_{\boldsymbol{\Phi}_{K}}}],$ (6)
@@
Figure 6: Trajectory smoothing of fragments leads to correction of Gaussians and help render better novel views.
@@
$\boldsymbol{\Delta}^{\textrm{merge}}{\boldsymbol{\Phi}_{ij}}=\lambda\boldsymbol{\Delta}{\boldsymbol{\Phi}_{i}}+(1-\lambda)\boldsymbol{\Delta^{*}}{\boldsymbol{\Phi}_{j}},$ (7)
@@
where $\lambda=0.5$, and only $\boldsymbol{\Delta^{*}}{\boldsymbol{\Phi}_{j}}$ is learnable. With intra-frame motions already smooth, we freeze $\boldsymbol{\Delta}{\boldsymbol{\Phi}_{i}}$ and $\boldsymbol{\Delta}{\boldsymbol{\Phi}_{j}}$, optimizing only $\boldsymbol{\Delta}^{\textrm{merge}}{\boldsymbol{\Phi}_{ij}}$ for a few iterations (e.g., 1,000) at a low learning rate using Eq.5 to ensure smooth inter-frame motion between fragments. Starting with $\boldsymbol{\Delta}{\boldsymbol{\Phi}_{1}}$, we progressively merge all deformation fields in a bottom-up fashion, resulting in a smooth and globally coherent 3D motionwithout abrupt transitions or flickers.
@@
Cascaded trajectory smoothing. Despite smooth transitions across fragments, minor 3D inconsistencies may persist (see Fig. 6), often due to disoriented Gaussians causing blurry or over-reconstructed artifacts [14]. Since the 3D Gaussian geometry is controlled by the covariance matrix (i.e., rotation $q$ and scaling $s$), we regularize $q$ and $s$ over a fixed window with neighboring frame constraints. We adopt off-the-shelf tracking (e.g., CoTracker [12]) and depth models (e.g., DepthAnything [45]) in a sliding window manner for post-processing. Given a window $w$, we estimate depth $\mathcal{D}$ and trajectory $\mathcal{T}$ on the video $\mathbf{V}$ (Eq. 2) and lift $N$ randomly selected trajectories to 3D using camera intrinsics. Points visible for at least $80\%$ frames are smoothed with Exponential Moving Average (EMA) on rotation and scale:
$\begin{split}q_{t}&=\frac{sin(\alpha.\theta)}{sin\theta}q_{t}+\frac{sin((1-\alpha).\theta)}{sin\theta}q_{t-1},\\ s_{t}&=\alpha s_{t}+(1-\alpha)s_{t-1},\end{split}$ (8)
@@
where $\theta$ is the angle between consecutive rotations, and $\alpha$ is the EMA decay factor. The process is repeated iteratively until all disoriented Gaussians are corrected, yielding stable and flicker-free 3D reconstructions.
4 Experiments and Results
@@
Dataset. We evaluate our method on 4D sequences with large object motions, defined when multiple object parts move. We introduce the I4D-15 benchmark, comprising 15 articulated objects across categories like Vehicles, Flowers, Humans, Animals, and Daily Life scenes. The dataset includes 64-frame sequences at 16 fps from Objaverse1.0 [7], rendered from 5 evenly spaced views at $0^{\circ}$ elevation. We select the first and last frames of the front view as input states and reserve the remaining video for evaluation. The camera radius is $1.5$, and the FOV is $49.1^{\circ}$, consistent with [11]. Motion filtering [19] ensures large motion selection. Evaluation uses appearance metrics (LPIPS, FVD)[30] and geometry metrics (SI-CD, CD)[27]. More details are provided in the supplementary.
@@
Figure 7: Effect of Inter-Fragment Consistency. Without using any consistency or regularization, blurriness and oversaturated artifacts are produced. Rigid consistency improves the structure and when combined with temporal-aware multi-view generation, better geometry and texture are obtained.
@@
Baselines. As we introduce a novel task, no existing method can be directly applied to our setting. Thus, we establish the following baselines for quantitative comparison: For 4D baseline generation, we first generate videos from two frames and subsequently lift it to 4D using a Video-to-4D approach. (a) Baseline-I employs a image based interpolation DreamMover [32] followed by [43] to generate video fragments and then use a latest Video-to-4D method [31] (b) Baseline-II utilizes latest video interpolation approach [43] along with our temporal hierarchy module for 2D interpolation and a recent Video-to-4D method [42]. Additionally, we evaluate a single-image-to-4D task on our dataset, with further analysis provided in the supplementary. Baseline results are obtained using official GitHub implementations.
@@
Quantitative comparisons. We quantitatively evaluate our approach on our I4D-15 benchmark. Two images from one view are used as input and 4 videos (each 64 frames) from other viewpoints and their corresponding timsetep point clouds are used for evaluation. As shown in Tab. 1, our method outperforms the baseline across all metrics in appearance and geometry. This shows the effectiveness of our method in handling complex motions in 4D by dividing it into quasi-static temporal fragments.
Table 1: Quantitative Analysis on proposed I4D-15 Dataset.
@@
Ours 0.91 0.103 679.23 22.67 0.59
Qualitative comparisons. Fig. 10 provides some qualitative comparisons with the baseline. It is apparent that our method produces less artifacts. Additional visual results are shown in Fig. 9 for different motions and object categories. More visual results are provided in the supplementary files.
Ablation Study. This study evaluates the contribution of key components in our method on the I4D-15 benchmark. As shown in Tab 2, using four segments enables our model to decompose complex motion into finer details, achieving the best cost-performance balance. Additionally, we visually analyze the effect of Intra-Fragment consistency (Sec. 3.3) in Fig. 7, revealing that mv-consistency significantly enhances novel view synthesis, while rigid consistency mitigates deformation artifacts. Tab. 3 further highlights the impact of 3D motion aggregation. The combination of merging and smoothing improves both appearance and geometry metrics, except for a slight decline in FVD when trajectory smoothing is applied. Moreover, we benchmark runtime against all baselines, as shown in Tab. 5, demonstrating that our approach outperforms the fastest baseline (B-I) by 25% on an NVIDIA A100 GPU. Additional ablations are provided in the supplementary material.
Table 2: Ablation on the impact of temporal fragment generation.
@@
4 (Ours) 0.103 679.23 22.67 0.59 17 mins
Table 3: Ablation on the components of motion aggregation.
@@
User study. A user study was conducted, as human judgment is most effective for assessing 3D generation and motion quality. We gathered 15 generated 4D motions and asked 20 participants to rank four methods (1 = best, 4 = worst) based on 3D geometry and motion consistency (reduced flicker). In case of ties in motion consistency, 3D generation quality was prioritized. As shown in Tab. 5, our method was preferred for overall 4D generation quality.
Table 4: Ablation on runtime
@@
Ours 679 35 min
Table 5: User study
@@
B-I 2.93
B-II 2.44
Ours 1.29
Application: Customized 4D Motion. In contrast to most existing 4D generation methods [30, 48] that depend on SDS [28], our approach improves controllability and motion diversity. While BLIP [17] is used by default to extract motion prompts, users can input custom prompts to generate 4D motions for the same initial and final states. As shown in Fig. 8, both jumping and walking motions of a dog are synthesized under identical start and end conditions. Despite motion complexity, our bottom-up 3D optimization ensures artifact-free novel view generation.
Application: customized 4D motion. In contrast to most existing 4D generation methods [30, 48] that depend on SDS [28], our approach improves controllability and motion diversity. While BLIP [17] is used by default to extract motion prompts, users can input custom prompts to generate 4D motions for the same initial and final states. As shown in Fig. 8, both jumping and walking motions of a dog are synthesized under identical start and end conditions. Despite motion complexity, our bottom-up 3D optimization ensures artifact-free novel view generation.
Figure 8: Controllable Motions. In-2-4D allows generation of diverse motions for the same start and end states
5 Conclusion, Limitations, Future Work

### v3

@@ 3.2. Temporal Fragment Hierarchy
where $\texttt{CS}(.)$ represents cosine similarity, and $p,q$ denote tokens of DIFT feature $f$. A frame is marked as a keyframe if the mean heatmap value of $H_{i,j}$ between frame pairs $I_{i},I_{j}$ falls below a predefined threshold. To sample the best keyframe in terms of visual fidelity we further assess its consistency with the initial inputs using FID metric. The keyframe latent $z_{m}$ at timestep $m$ is selected based on the highest FID against the input states to remain faithful to inputs. For instance, in Fig 4, the chosen keyframe exhibits the highest fidelity to the input states of the eagle. Once identified, the keyframe divides the motion trajectory into two segments: $z_{s},z_{m}$} and $z_{m},z_{e}$}. The interpolation model $\psi(.)$ then utilizes these fragments iteratively in a ”divide-and-conquer” fashion, identifying further keyframes until the full video is processed. This hierarchical approach ensures adaptive keyframe density, reducing redundant intermediate frames in low-motion areas while preserving complex motion details. Therefore, the hierarchical keyframe selection is performed recursively based on prior selections:
(1) $\mathcal{K}={K_{(s)(1)},K_{(1)(2)},K_{(2)(3)},...,K_{(c)(e)}}$
@@
(2) $\mathbf{V}=\{\mathcal{V}_{s(1f)},\mathcal{V}_{(1f)(2f)},...,\mathcal{V}_{((c-1)f)e}\},$
@@
3.3. Modelling Intra-Fragment Geometry
We lift individual video fragments to 4D by generating multi-view videos of the object. Existing video-to-4D methods (Ren et al., 2023; Wu et al., 2024) use multi-view diffusion models (Liu et al., 2023) to synthesize multi-view videos by independently processing each frame. However, this approach ensures cross-view consistency but leads to temporally inconsistent geometry Fig 7. Moreover, due to sole reliance on multi-view video supervision, Gaussian splatting often produces flickering and texture variations (Luiten et al., 2024) due to its high degrees of freedom per point and lack of motion constraints. We address these issues by generating temporally consistent multi-view videos and regularizing motion with rigid constraints within each fragment.
Learning canonical 3D. We start by estimating a canonical Gaussian representation and then add motion to it from the multi-view videos. For each temporal fragment $\mathcal{V}_{i}$, we designate the keyframe $K_{(i)(j)}$ as the canonical reference and reconstruct its 3D structure via multi-view synthesis. Specifically, we employ the multiview diffusion model Era3D (Li et al., 2024a) to generate multi-view images from $K_{(i)(j)}$, followed by 3DGS (Kerbl et al., 2023a) for coarse geometry reconstruction. As each fragment is processed independently, parallel execution reduces computation time. The resulting coarse geometry provides an effective initialization for learning texture and geometry across the remaining temporal frames.
Dynamic 3D fragment generation. After learning 3D static Gaussians, we leverage motion priors from the video fragment to transform them into dynamic Gaussians. Since single-view videos cannot provide diverse observations of the scene from different viewpoints, we use multi-view videos. To promote temporal consistency, rather than generating multiple-view of the frames independently at each timestep, we propagate the self-attention features of the multi-view diffusion model (Li et al., 2024a) from the canonical frame across the entire frames of the fragment as follows:
@@ 3.3. Modelling Intra-Fragment Geometry
where ${z}_{c}$ is the multi-view latent of the canonical frame, $t$ is timestep of the video fragment, $\gamma$ is the blending weight and $d_{k}$ is the key dimension. With quasi-static motions in each video fragment, the generated multi-view videos have minimal variation in viewpoints, making it easier for the model to capture accurate and consistent geometry (Fig. 7). With the synthesized multi-view videos of the dynamic object, we optimize a 3D deformation field (denoted by $\Delta_{\boldsymbol{\Phi}_{i}}$) to enable free-viewpoint rendering. We chose Hexplanes (Cao and Johnson, 2023) as our deformation field due to modeling efficiency. The deformation field predicts each Gaussian’s geometric offsets at a given timestamp relative to the mean canonical state (keyframe). For each timestamp $\tau$ of video and 3D Gaussian $p$, Hexplanes predict displacement, rotation, and scaling for the 3D gaussian points.
Figure 5. Effect of inbetweening on geometry. When the input states are significantly different, the 3D deformation module undergoes large movements (fast motion) leading to artifacts in novel views, whereas generating intermediate frames between the states (slow motion) enhances the geometry using smaller deformations.
@@
[…]
where $I^{\tau}_{\textrm{Ref}}$ is the $\tau$-th frame, $o_{\textrm{Ref}}$ is the reference viewpoint, and $f$ is the rendering function. Dynamic Gaussians tend to move freely across regions of similar color (Luiten et al., 2024) without constraints, causing flickering and floating artifacts that degrade 4D motion realism. As motion within each fragment is minimal, we enforce rigid assumptions on point movements relative to the canonical state by regularization:
[…]
(5) $\mathcal{L}=\lambda_{1}\mathcal{L}_{Ref}+\lambda_{2}\mathcal{L}_{mask}+\lambda_{3}\mathcal{L}_{rigid},$
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are loss weights which are set as $0.4,0.3,0.3$ respectively. In training, for each fragment $\mathcal{V}_{i}$, we first use $\mathcal{L}$ to supervise the static 3D Gaussian, then train the dynamic 4D Gaussian with all reference frames.
3.4. Cascaded 3D Motion Aggregation
@@
(6) $\Delta=[\boldsymbol{\Delta_{\boldsymbol{\Phi}_{1}}},\boldsymbol{\Delta_{\boldsymbol{\Phi}_{2}}},...,\boldsymbol{\Delta_{\boldsymbol{\Phi}_{K}}}],$
@@
Figure 6. Trajectory smoothing of fragments leads to correction of Gaussians and help render better novel views.
@@
(7) $\boldsymbol{\Delta}^{\textrm{merge}}{\boldsymbol{\Phi}_{ij}}=\lambda\boldsymbol{\Delta}{\boldsymbol{\Phi}_{i}}+(1-\lambda)\boldsymbol{\Delta^{*}}{\boldsymbol{\Phi}_{j}},$
@@
where $\lambda=0.5$, and only $\boldsymbol{\Delta^{*}}{\boldsymbol{\Phi}_{j}}$ is learnable. With intra-frame motions already smooth, we freeze $\boldsymbol{\Delta}{\boldsymbol{\Phi}_{i}}$ and $\boldsymbol{\Delta}{\boldsymbol{\Phi}_{j}}$, optimizing only $\boldsymbol{\Delta}^{\textrm{merge}}{\boldsymbol{\Phi}_{ij}}$ for a few iterations (e.g., 1,000) at a low learning rate using Eq.5 to ensure smooth inter-frame motion between fragments. Starting with $\boldsymbol{\Delta}{\boldsymbol{\Phi}_{1}}$, we progressively merge all deformation fields in a cumulative fashion, resulting in a smooth and globally coherent 3D motionwithout abrupt transitions or flickers.
@@
Cascaded trajectory smoothing. Despite smooth transitions across fragments, minor 3D inconsistencies may persist (see 6), often due to disoriented Gaussians causing blurry or over-reconstructed artifacts (Kerbl et al., 2023b). Since the 3D Gaussian geometry is controlled by the covariance matrix (i.e., rotation $q$ and scaling $s$), we regularize $q$ and $s$ over a fixed window with neighboring frame constraints. We adopt off-the-shelf tracking (e.g., CoTracker (Karaev et al., 2023)) and depth models (e.g., DepthAnything (Yang et al., 2024a)) in a sliding window manner for post-processing. Given a window $w$, we estimate depth $\mathcal{D}$ and trajectory $\mathcal{T}$ on the video $\mathbf{V}$ (Eq. 2) and lift $N$ randomly selected trajectories to 3D using camera intrinsics. Points visible for at least $80\%$ frames are smoothed with Exponential Moving Average (EMA) on rotation and scale:
(8) $\begin{split}q_{t}&=\frac{sin(\alpha.\theta)}{sin\theta}q_{t}+\frac{sin((1-\alpha).\theta)}{sin\theta}q_{t-1},\\ s_{t}&=\alpha s_{t}+(1-\alpha)s_{t-1},\end{split}$
@@
where $\theta$ is the angle between consecutive rotations, and $\alpha$ is the EMA decay factor (set as 0.6). The process is repeated iteratively until all disoriented Gaussians are corrected, yielding stable and flicker-free 3D reconstructions.
4. Experiments and Results
@@
Dataset. We evaluate our method on 4D sequences with large object motions, defined when multiple object parts move. We introduce the I4D-15 benchmark, comprising 15 articulated objects across categories like Vehicles, Flowers, Humans, Animals, and Daily Life scenes. The dataset includes 64-frame sequences at 16 fps from Objaverse1.0 (Deitke et al., 2023) and SketchFab (Spiess et al., 2024), rendered from 5 evenly spaced views at $0^{\circ}$ elevation. We select the first and last frames of the front view as input states and reserve the remaining video for evaluation. The camera radius is $1.5$, and the FOV is $49.1^{\circ}$, consistent with (Jiang et al., 2023). Motion filtering (Li et al., 2024b) ensures large motion selection. Evaluation uses appearance metrics (LPIPS, FVD)(Ren et al., 2023) and geometry metrics (SI-CD, CD)(Peng et al., 2024).In addition to this, we have also evaluated our method along with the baselines on Consistent-4D (Jiang et al., 2023) dataset for more extensive validation.
@@
Figure 7. Effect of Inter-Fragment Consistency. Without using any consistency or regularization, blurriness and oversaturated artifacts are produced. Rigid consistency improves the structure and when combined with temporal-aware multi-view generation, better geometry and texture are obtained.
Figure 8. Qualitative comparison of our approach with the baselines on challenging synthetic and real-world scenes (e.g, water splash , weightlifting).
@@
Baselines. As we introduce a novel task, no existing method can be directly applied to our setting. Thus, we establish the following baselines for quantitative comparison: For 4D baseline generation, we first generate videos from two frames and subsequently lift it to 4D using a Video-to-4D approach. (a) Baseline-I employs a image based interpolation DreamMover (Shen et al., 2024) to generate keyframes followed by (Xing et al., 2024) to generate video fragments and then use a latest Video-to-4D method (Ren et al., 2025) (b) Baseline-II utilizes latest video interpolation approach (Xing et al., 2024) along with our temporal fragment generation module for 2D interpolation and a recent Video-to-4D method (Xie et al., 2024). (c) Baseline-III is an improved version of Baseline I in two ways: first, it incorporates the output of our Temporal Fragment Hierarchy Module (Sec. 3.2) for video generation and second, we replace L4GM (Ren et al., 2025) with a more recent video-to-4D model, SV4D (Xie et al., 2024) to lift the video to 4D. Additionally, we evaluate a single-image-to-4D task on our dataset, with further analysis provided in the supplementary. Baseline results are obtained using official GitHub implementations.
@@
Quantitative comparisons. We quantitatively evaluate our approach on our I4D-15 benchmark. Two images from one view are used as input and 4 videos (each 64 frames) from other viewpoints and their corresponding timsetep point clouds are used for evaluation. As shown in Tab. 1, our method outperforms the baseline across all metrics in appearance and geometry. To compare our method on the Consistent4D benchmark, we use the first and the last frame of the input test video and use the novel view videos of the test set for evaluation. Similar observations can be drawn from Tab. 2, where our method is much superior than the baselines which is expected as the dataset contains simple motions. This shows the robustness of our model design and effectiveness of our method in handling complex motions in 4D by dividing it into quasi-static temporal fragments.
Table 1. Quantitative Analysis on proposed I4D-15 Dataset.
@@
Baseline-III 0.83 0.138 811.26 32.46 0.74
[…]
Table 2. Quantitative Analysis on proposed Consistent-4D Dataset.
Method LPIPS $\downarrow$ CLIP-S $\uparrow$ FVD $\downarrow$
Baseline-I 0.18 0.84 925.46
Baseline-II 0.14 0.87 810.24
Baseline-III 0.16 0.86 878.13
[…]
Qualitative comparisons. Fig. 8 and Fig  11 provides some qualitative comparisons with the baseline on both real-world scenes and synthetic objects. It is apparent that our method produces fewer artifacts in the novel views and also preserve the geometry better in complex motions. In addition to this, we also present a gallery visualization of scenes in Fig 10 for different motions and object categories from our proposed I4D-15 benchmark and additional real-world scenes. More videos are provided in supplementary.
Ablation Study. This study evaluates the contribution of key components in our method on the I4D-15 benchmark. As shown in Tab 3, using four segments enables our model to decompose complex motion into finer details using our Temporal Fragment Hierarchy (TFH) module, achieving the best cost-performance balance. Additionally, we visually analyze the effect of Intra-Fragment consistency (Sec. 3.3) in Fig. 7, revealing that mv-consistency significantly enhances novel view synthesis, while rigid consistency mitigates deformation artifacts. Tab. 4 further highlights the impact of 3D motion aggregation. The combination of merging and smoothing improves both appearance and geometry metrics, except for a slight decline in FVD when trajectory smoothing is applied. Moreover, we benchmark runtime against all baselines, as shown in Tab. 6, demonstrating that our approach outperforms the fastest baseline (B-I) by 25% on an NVIDIA A100 GPU. Additional ablations are provided in the supplementary material.
Table 3. Ablation on the impact of temporal fragment generation.
@@
[…]
Table 4. Ablation on the components of motion aggregation.
@@
User study. A user study was conducted, as human judgment is most effective for assessing 3D generation and motion quality. Since no standard metrics exist for evaluating 4D performance holistically, human judgment becomes a viable option. The study involved $20$ graduate students with a background in computer science, who were shown both the start/end state images and the output rendered videos from all the methods. Each object motion was presented from four canonical views (front, back, left, and right) and the participants were asked to rank four methods (1 = best, 4 = worst) based on how well the start/end state images align with the generated 3D geometry and how consistent is the motion (reduced flicker). In case of ties in motion consistency, 3D generation quality was prioritized. As shown in Tab. 6, our method achieves the best average rank (1.46), demonstrating the superiority of our model for overall 4D generation quality.
Table 5. Ablation on runtime
@@
B-III 811 1.20 hr
[…]
Table 6. User study
@@
B-I 3.14
B-II 2.35
B-III 2.85
[…]
Application: Customized 4D Motion. In contrast to most existing 4D generation methods (Ren et al., 2023; Zeng et al., 2024) that depend on SDS (Poole et al., 2022), our approach improves controllability and motion diversity. While BLIP (Li et al., 2022) is used by default to extract motion prompts, users can input custom prompts to generate 4D motions for the same initial and final states. As shown in Fig. 9, both jumping and walking motions of a dog are synthesized under identical start and end conditions. Despite motion complexity, our bottom-up 3D optimization ensures artifact-free novel view generation.
Figure 9. Controllable Motions. In-2-4D allows generation of diverse motions for the same start and end states
5. Conclusion, Limitations, Future Work